Grade 11Physical Chemistry

what is the wavelength of light emitted when the electron in a hydrogen atom undergoes transition from an energy level with n=4 to an energy level with n=2?

Rishi Manoj
8 Years agoGrade 11

2 Answers

Arun
8 Years ago
Dear student
 
ṽ = R(1/n1 1/n22) = 109677(1/2 1/42) cm-1 = 20564.4 cm-1 
λ = 1/ν = 1/20564.4 = 486×10-7 cm =  486×10-9 m = 486 nm
 
 
Regards
Arun (askIITians forum expert)
Khimraj
8 Years ago
Wave number is given by
ṽ = R(1/n1 1/n22) = 109677(1/2 1/42) cm-1 = 20564.4 cm-1 
λ = 1/ν = 1/20564.4 = 486×10-7 cm =  486×10-9 m = 486 nm
hope it clears.